OpenStudy (anonymous):

Noah currently has an account balance of $2,976.45. He opened the account 11 years ago with a deposit of $1,963.45. If the interest compounds quarterly, what is the interest rate on the account?

OpenStudy (anonymous):

The formula you need for this one is\[A(t)=P(1+\frac{r}{n})^{nt}\]A(t) is the amount in the account. P is the pricipal (amount deposited). n is the number of times per year the account is compounded. r is the inereest rate. t is time (in years)

OpenStudy (anonymous):

Setting up the particulars of this problem: \[2976.45=1963.45(1+\frac{r}{4})^{4(11)}\]

OpenStudy (anonymous):

Can you solve that?